OpenStudy (anonymous):

Which is a factor of 8x2 – 12x – 8 ? a.) x + 1 b.) x – 4 c.) 2x – 2 d.) 2x + 1

OpenStudy (anonymous):

use the factor theorem eg if x+1 is a factor then f(-1) = 0 if 2x +1 is a factor then f(-1/2) = 0 try x+1; f(-1) = 8(-1)^2 - 12(-1) - 8 = 8+12 -8 = 12 - so x+1 is not a factor

OpenStudy (anonymous):

or you can try factorising 8x^2 - 12x - 8 =4(2x^2 - 3x - 2) = 4(2x + 1)(x - 2) - thats it

OpenStudy (callisto):

I suggest you to do factorization 8x^2 – 12x – 8 = 4 (2x^2 -3x -2) = 4 ( 2x +1) (x-2)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Got it... 8x2-12x-8=4(2x2-3x-2) =4(2x2+x-4x-2) =4(x(2x+1)-2(2x+1)) =4(x-2)(2x+1)